### v1

# FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs

Abstract
Large Language Models (LLMs) frequently generate hallucinated content, posing significant challenges for applications where factuality is crucial. While existing hallucination detection methods typically operate at the sentence level or passage level, we propose FactSelfCheck, a novel black-box sampling-based method that enables fine-grained fact-level detection. Our approach represents text as knowledge graphs consisting of facts in the form of triples. Through analyzing factual consistency across multiple LLM responses, we compute fine-grained hallucination scores without requiring external resources or training data. Our evaluation demonstrates that FactSelfCheck performs competitively with leading sampling-based methods while providing more detailed insights. Most notably, our fact-level approach significantly improves hallucination correction, achieving a $35\%$ increase in factual content compared to the baseline, while sentence-level SelfCheckGPT yields only an 8% improvement. The granular nature of our detection enables more precise identification and correction of hallucinated content.
FactSelfCheck: Fact-Level Black-Box Hallucination Detection for LLMs
Albert Sawczyn1                        Jakub Binkowski1                        Denis Janiak1
Bogdan Gabrys2                        Tomasz Kajdanowicz1
1Wrocław University of Science and Technology 2University of Technology Sydney albert.sawczyn@pwr.edu.pl
1 Introduction
Large Language Models (LLMs) have gained significant attention from academia and industry recently. However, a major limitation of LLMs is their tendency to generate hallucinated information Farquhar et al. (2024); Huang et al. (2025), posing significant challenges for applications where factual correctness is crucial, such as healthcare Sallam (2023). Although numerous methods have been proposed to reduce hallucinations Zhang et al. (2023), it is not possible to eliminate them, and LLMs will constantly hallucinate Lee (2023); Xu et al. (2024). Therefore, there remains a critical need for reliable hallucination detection in LLM responses. Effective detection enables system interventions by either preventing the transmission of hallucinated content to users or facilitating its correction Zhang et al. (2023).
Previous approaches to hallucination detection have primarily focused on classifying hallucinations at either the passage or sentence level Huang et al. (2025). While valuable, these approaches are limited in their granularity, as they do not provide detailed information about specific hallucinated facts. To address this limitation, we propose a novel method for hallucination detection that operates at the fact level, offering finer-grained analysis. In our approach, we define a fact as a triple consisting of a head, relation, and tail – a standard representation in knowledge graphs (e.g., (Robert Smithmember ofThe Cure)) Hamilton et al. (2017). Our method provides more precise and actionable information by computing hallucination scores for individual facts than traditional passage-level or sentence-level classification approaches. Moreover, our method represents facts in a structured way through knowledge graphs (KGs), enabling more straightforward interpretation, processing, and verification of factual content Pan et al. (2024).
Figure 1: The pipeline of FactSelfCheck in two variants. For response $p$, entities $\mathcal{E}_{p}$ and relations $\mathcal{R}_{p}$ are extracted, followed by the construction of knowledge graphs $KG_{p}$, for which hallucination scores $\mathcal{H}_{\text{fact}}$ are calculated. Samples’ entities $\mathcal{E}_{S}$ and relations $\mathcal{R}_{S}$ are created by merging $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$ with entities and relations from $KG_{p}$. For each sample $s$, the knowledge graph $KG_{s}$ is extracted. FactSelfCheck-KG assesses the consistency between a fact and all $KG_{s}$. FactSelfCheck-Text assesses the consistency between a fact and all $s$ directly. To obtain sentence-level $\mathcal{H}_{\text{sentence}}$ and passage-level $\mathcal{H}_{\text{passage}}$ scores, fact-level scores are aggregated, as indicated by dashed arrows.
Our granular approach is motivated by two key observations. First, a single sentence or text passage can contain multiple facts, with the number of facts varying significantly across sentences, contexts, and domains. This variability makes it challenging to identify hallucinated aspects of generated output precisely when using sentence-level or passage-level detection. Second, false information can be dispersed throughout a text, as a single fact may appear across multiple sentences. That can mislead the sentence-level detection, as the factuality of a sentence is dependent on the previous sentences Zhang et al. (2024). These challenges become particularly acute when analyzing long texts generated by LLMs. Fine-grained fact-level detection provides a more precise understanding of text factuality than a sentence or passage analysis. It enables better assessment of content reliability and, as we show later, more effective factuality correction.
We propose FactSelfCheck, a black-box method for fact-level hallucination detection, meaning it does not require access to the model’s internal parameters. This design choice makes our approach universally applicable across any LLM, including closed models, like GPT OpenAI et al. (2024). Following a sampling-based detection paradigm, introduced by Manakul et al., 2023, our method utilizes multiple response generations and analyzes the factual consistency of extracted facts across these samples. This paradigm is based on the phenomenon that factual information remains largely consistent across different generations, while hallucinated content tends to vary or contradict itself between samples Manakul et al. (2023); Wang et al. (2023). This way, we can effectively identify hallucinated facts without relying on external resources (zero-resource) or access to the model’s internal parameters (black-box). Moreover, our method is non-parametric, as it does not require any training, making it easy to implement in any domain without the need for training data. The FactSelfCheck pipeline consists of three main steps: knowledge graph extraction, which extracts sets of facts from the initial response and samples; fact-level hallucination scoring; and calculating sentence-level and passage-level scores by aggregating fact-level scores.
We evaluated our method using the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023), aggregating fact-level scores into sentence-level and passage-level scores. Our approach achieves performance comparable to leading sampling-based methods while providing more detailed information about hallucinations. Additionally, we demonstrate that our fact-level approach significantly improves hallucination correction. Compared to a baseline, providing incorrect facts to the correction method leads to a $35\%$ increase in factual content, while passing incorrect sentences leads to an $8\%$ increase.
Our key contributions are as follows:
  1. 1.
    The novel black-box sampling-based method for fact-level hallucination detection – FactSelfCheck. It enables fine-grained hallucination detection in LLM responses without requiring training data or external resources, as it is both non-parametric and zero-resource.
  2. 2.
    The effective approaches for measuring factual consistency across multiple samples: FactSelfCheck-KG using knowledge graph comparisons and FactSelfCheck-Text using direct text comparison.
  3. 3.
    Comprehensive evaluation of our method, which shows competitive performance with leading sampling-based methods while providing more detailed insights.
  4. 4.
    Demonstration that fact-level detection significantly improves hallucination correction compared to sentence-level approaches.
Our code is available on GitHub  [^1] https://github.com/graphml-lab-pwr/FactSelfCheck license: CC BY-SA 4.0.
2 Related work
Xu et al., 2024 have proven that hallucinations are inevitable in LLMs. As LLMs are powerful tools, many recent studies have been conducted regarding hallucination mitigation and detection Zhang et al. (2023); Huang et al. (2025). The detection methods can be divided into two groups: white-box and black-box.
White-box methods analyze LLMs’ internal states (Farquhar et al., 2024; Azaria and Mitchell, 2023). While these methods are universal across all LLMs, they often require multiple generations, similar to sampling-based methods. Notable approaches include: SAPLMA (Azaria and Mitchell, 2023), which predicts from hidden states whether generated text is correct or incorrect; INSIDE (Chen et al., 2024), which evaluates hidden state consistency across generations; SEPs (Kossen et al., 2024) that predict entropy directly from model hidden states; Lookback Lens (Chuang et al., 2024) and AttentionScore (Sriramanan et al., 2024) that uses attention maps to detect hallucinations.
Black-box approaches operate without access to the model’s internal states and aim to detect hallucinations based solely on the text generated by LLMs. Some of these methods use external resources to collect evidence (Min et al., 2023; Chern et al., 2023). Others leverage LLMs to detect hallucinations like CoVe Dhuliawala et al. (2024), which utilizes the chain-of-thought paradigm for detection. Another category is sampling-based methods, such as SelfCheckGPT (Manakul et al., 2023), which evaluate factuality by generating multiple responses (stochastic samples) and assessing consistency between the original response and these samples.
The most popular approach is to classify hallucinations at sentence-level or passage-level Huang et al. (2025). Few methods have been specifically designed to detect hallucinations at the fact level (where facts are defined as triples). GraphEval Sansford et al. (2024) generates a KG from LLM output and compares it with the context provided in the LLM input. FactAlign Rashad et al. (2024) builds KGs from input and output, then compares them after performing entity alignment, a technique that pairs the same entity in different KGs. Most similar to our approach is GCA Fang et al. (2024), which constructs KGs from the response and samples and then compares them by aggregating multiple scores. GCA has significant methodological concerns - they tuned several parameters directly on the evaluation set (which is the only set in this dataset), thus we do not compare against it despite using the same dataset. Moreover, our method belongs to a fundamentally different class as it is zero-shot, requiring no parameter tuning.
3 Method
We propose FactSelfCheck, a black-box sampling-based method for fact-level hallucination detection, as illustrated in Figure 1. We list used prompts in Appendix F.
3.1 Notation
Let $p$ denote the initial response passage generated by the LLM to a user query, which we aim to evaluate for hallucinations. Let $S=\{s_{1},\ldots,s_{N}\}$ represent a set of $N$ stochastic LLM response samples. The text passage $p$ consists of a set of sentences $U$. For each sentence $u\in U$ and each sample $s\in S$, we extract knowledge graphs $KG_{u}$ and $KG_{s}$, respectively. Each knowledge graph comprises a set of facts, where a fact $f$ is defined as a triple $(h,r,t)$ consisting of a head $h$, relation $r$, and tail $t$, e.g. (Robert Smith, member of, The Cure). We define $KG_{p}=\bigcup_{u\in U}KG_{u}$ as the knowledge graph consisting of all facts from the passage $p$.
Our objective is to compute a fact-level hallucination score $\mathcal{H}_{\text{fact}}$ for each fact $f$ in $KG_{p}$. Subsequently, to facilitate comparisons with other methods, we aggregate these scores to obtain a sentence-level hallucination score $\mathcal{H}_{\text{sentence}}$ for each sentence $u$ and a passage-level hallucination score $\mathcal{H}_{\text{passage}}$ for the entire passage $p$.
3.2 FactSelfCheck pipeline
As shown in Figure 1, the pipeline of FactSelfCheck consists of three main steps: (1) Knowledge Graph Extraction that extracts sets of entities, relations, and finally, knowledge graph from the initial response $p$ and samples $S$; (2) Fact-level Hallucination Scores, that score facts by measuring factual consistency between facts in $KG_{p}$ and, depending on the variant, $KG_{s}$ in FactSelfCheck-KG or directly $s$ in FactSelfCheck-Text; (3) Sentence-Level and Passage-Level Scores calculation by aggregation of the fact-level scores.
3.3 Knowledge Graph Extraction
We adopt an approach that decomposes the knowledge graph extraction task into simpler subtasks, similarly to Edge et al., 2024. This process involves three primary steps: extracting entities, identifying relations, and formulating facts, which are implemented as a sequence of LLM prompts.
For each instance, we extract a list of entities $\mathcal{E}_{p}$ by passing $p$ to the LLM (Prompt 1).
$\mathcal{E}_{p}=LLM_{\text{entities}}(p)$ (1)
Next, we provide $p$ and $\mathcal{E}_{p}$ to the LLM to extract relation types between entities, resulting in $\mathcal{R}_{p}$ (Prompt 2).
$\mathcal{R}_{p}=LLM_{\text{relations}}(p,\mathcal{E}_{p})$ (2)
We then input $\mathcal{E}_{p}$, $\mathcal{R}_{p}$, and each sentence $u\in U$ into the LLM to extract the knowledge graph $KG_{u}$ (Prompt 3). We also provide an initial response $p$ to add contextual information. The output is a set of facts:
$KG_{u}=LLM_{\text{sentence\_facts}}(u,p,\mathcal{E}_{p},\mathcal{R}_{p})$ (3)
After extracting $KG_{u}$ for each sentence $u\in U$, we compile the sets of entities $\mathcal{E}_{S}$ and relations $\mathcal{R}_{S}$ required for extracting knowledge graph $KG_{s}$ from each sample $s$ (Prompt 4) [^2] Although we could theoretically use $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$ for $KG_{s}$ extraction, in practice, LLMs are not sufficiently accurate to extract all entities and relations from the response $p$ when calculating $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$. This results in $KG_{u}$ containing entities and relations not present in $\mathcal{E}_{p}$ and $\mathcal{R}_{p}$, even if the prompt restricts them. Empirical tests showed that extending $\mathcal{E}_{S}$ and $\mathcal{R}_{S}$ by adding entities and relations from all $KG_{u}$ improved the results.:
$\mathcal{E}_{S}=\mathcal{E}_{p}\cup\bigcup_{u\in U}\{h,t\mid(h,r,t)\in KG_{u}\}$ (4)
$\mathcal{R}_{S}=\mathcal{R}_{p}\cup\bigcup_{u\in U}\{r\mid(h,r,t)\in KG_{u}\}$ (5)
$KG_{s}=LLM_{\text{sample\_facts}}(s,\mathcal{E}_{S},\mathcal{R}_{S})$ (6)
Extracting $KG_{s}$ by utilizing $\mathcal{E}_{S}$ and $\mathcal{R}_{S}$ is more convenient than extraction without them, as it eliminates the need for entity alignment and ensures that the KG is built using the same schema as $KG_{p}$.
3.4 Fact-Level Hallucination Scores
We define two variants for measuring fact-level hallucination scores. The first variant, FactSelfCheck-KG, assesses the consistency between a fact and the knowledge graphs extracted from samples. The second variant, FactSelfCheck-Text, evaluates the consistency between a fact and the samples directly.
3.4.1 FactSelfCheck-KG
In the FactSelfCheck-KG variant, we introduce two metrics to assess the reliability of each fact.
Frequency-Based Hallucination Score
The frequency-based fact-level hallucination score is based on the intuition that the probability of a fact being hallucinated is inversely proportional to the fraction of samples containing the same fact.
$\mathcal{H}_{\text{fact}}(f)=1-\frac{1}{|S|}\sum_{s\in S}\mathbb{I}\{f\in KG_{s}\}$ (7)
where $\mathcal{H}_{\text{fact}}(f)$ is the hallucination score for fact $f$, and $\mathbb{I}\{f\in KG_{s_{n}}\}$ is an indicator function that equals $1$ if fact $f$ appears in $KG_{s_{n}}$ and 00 otherwise. The higher the $\mathcal{H}_{\text{fact}}$ value, the higher the plausibility of hallucination.
LLM-Based Hallucination Score
To allow semantic matching and reasoning over knowledge graphs, rather than only exact fact matching, we introduce the LLM-based fact-level hallucination score. We instruct the LLM to determine whether each fact is supported by the knowledge graphs extracted from the samples (Prompt 5). The LLM is expected to respond with ’yes’ or ’no’. We then average the valid responses from the LLM to get the final score as in Equation (8). Any invalid responses are not included in the averaging.
$\mathcal{H}_{\text{fact}}(f)=\frac{1}{|V_{f}|}\sum_{s\in V_{f}}\Psi(f,KG_{s})$ (8)
where $V_{f}$ represents the set of samples with valid LLM responses for the fact $f$, and the function $\Psi$ is defined as follows:
$\Psi(\cdot)=\begin{cases}0&\text{if the LLM returns 'yes'}\\ 1&\text{if the LLM returns 'no'}\end{cases}$ (9)
3.4.2 FactSelfCheck-Text
In the FactSelfCheck-Text variant, we check if a fact is supported by each textual sample directly without using the knowledge graphs. We prompt the LLM to evaluate whether a fact $f$ is supported by the textual sample $s$ (Prompt 5). As in the previous variant, we average the valid LLM responses using the $\Psi$ function:
$\mathcal{H}_{\text{fact}}(f)=\frac{1}{|V_{f}|}\sum_{s\in V_{f}}\Psi(f,s)$ (10)
3.5 Sentence-Level and Passage-Level Hallucination Scores
While detecting hallucinations at the fact level offers fine-grained insights, there are scenarios where sentence-level or passage-level detection is necessary. To achieve this, we aggregate fact-level scores to compute sentence-level scores $\mathcal{H}_{\text{sentence}}(u)$ and a passage-level score $\mathcal{H}_{\text{passage}}(p)$.
$\mathcal{H}_{\text{sentence}}(u)=\text{Agg}_{f\in KG_{u}}\mathcal{H}_{\text{fact}}(f)$ (11)
$\mathcal{H}_{\text{passage}}(p)=\text{Agg}_{u\in U}\mathcal{H}_{\text{sentence}}(u)$ (12)
where $u$ represents a single sentence, and $U$ is the set of sentences of the response $p$. The aggregation function Agg can be based on operations such as $mean$ or $max$. The $mean$ function provides a smoothed factuality score, whereas $max$ is sensitive to the most hallucinated facts or sentences. Additionally, sentences without facts are excluded from the aggregation when calculating the passage-level score.
As discussed in Section 4.1, the dataset choice influences the calculation of the passage-level score in Equation 12. Since the ground truth scores are derived by averaging sentence-level scores, we also employ the $mean$ function for aggregation. It is worth mentioning that a more intuitive approach would be to average the fact-level scores from $KG_{p}$, but such an approach cannot be evaluated with the current dataset. Nevertheless, this technique could be promising for real-world applications and future research.
4 Experimental Setup
In this section, we describe the experimental setup, including the used data, implementation details, and aspects we investigated.
4.1 Evaluation Data
We utilized the WikiBio GPT-3 Hallucination Dataset Manakul et al. (2023) 333huggingface.co/datasets/potsawee/wiki_bio_gpt3_hallucination (cc-by-sa-3.0 license) for all experiments. It is designed to evaluate sampling-based hallucination detection methods. It consists of $238$ passage-level annotations and $1908$ sentence-level annotations. As it contains only test data, tuning parameters on this dataset would not be methodologically correct. While the dataset focuses on sentence-level and passage-level evaluation, our approach goes further by providing more fine-grained insights. To ensure a meaningful comparison with SelfCheckGPT, we evaluated these levels using aggregation approaches (see Section 3.5). Following the protocol established by Manakul et al., 2023, we merged the labels major-inaccurate and minor-inaccurate into a single hallucinated class.
Additionally, we conducted a fact-level evaluation to provide a more comprehensive analysis, with the detailed description and results presented in Appendix B.
4.2 Implementation Details
We employed the Llama-3.1-70B-Instruct model Grattafiori et al. (2024) as the LLM in our method, hosting it locally using vLLM Kwon et al. (2023) on a server with 2xNvidia H100 94GB. The sole exception was the hallucination correction experiment, for which we utilized GPT-4o OpenAI et al. (2024), as detailed in Section 4.5. We set the LLM’s temperature to $0.0$ for all calls, except during hallucination correction (see Section 4.5), where we set it to $0.5$. We implemented the methods and experiments using LangChain Chase (2022) and Hugging Face Datasets Lhoest et al. (2021). All pipeline steps were defined using DVC Kuprieiev et al. (2025) to facilitate reproducibility.
To determine whether a language model’s response was ’yes’ or ’no’ in Equation 9, we parsed the text into individual words and verified the presence of the words ’yes’ or ’no’. If either word was detected, the response was excluded from the averaging process in Equations 8 and 10.
4.3 Evaluation of Sentence-Level and Passage-Level Hallucination Detection
For a fair comparison, we employed the same evaluation protocol as SelfCheckGPT. We reported area under the precision-recall curve (AUC-PR) for the sentence-level task and Pearson and Spearman correlation coefficients for the passage-level task. We evaluated our method against all variants of SelfCheckGPT, including Prompt, NLI, Unigram (Max), QA, and BERTScore. We reproduced the results of the Prompt variant using the same LLM employed in our study, namely Llama-3.1-70B-Instruct, whereas the original method utilized an unspecified release of GPT-3.5-turbo. We obtained the remaining results of SelfCheckGPT from the original paper. We ensured consistency in the evaluation protocol by reviewing their source code 444github.com/potsawee/selfcheckgpt.
4.4 Effect of Sample Size on Detection Performance
We investigated the impact of varying the number of samples on the detection performance of our method. Specifically, we evaluated the performance at both the sentence and passage levels by changing the number of samples from $1$ to $20$. We compared our method with the SelfCheckGPT (Prompt) approach.
4.5 Role of Fact-Level Detection in Hallucination Correction
One potential application of hallucination detection methods is their use in correcting hallucinated responses. In this experiment, we investigate the effectiveness of our fact-level detection approach in enhancing hallucination correction and compare the results with those obtained using sentence-level detection and a baseline method. Each of the three tested approaches uses different input for the LLM: (1) Baseline: the original prompt and the generated response (Prompt 7). (2) Sentence-level: the original prompt, the generated response, and a list of hallucinated sentences (Prompt 8). (3) Fact-level: the original prompt, the generated response, and a list of hallucinated facts (Prompt 9).
The original prompt is the one used during the creation of the dataset: "This is a Wikipedia passage about {concept_name}:". We instructed the LLM to return a list of sentences, allowing it to correct each sentence or leave it unchanged if no hallucinations were detected. We obtained the lists of incorrect sentences/facts using the best variants of models (see Section 5.1) with thresholds that achieved the highest F1-scores on the dataset ($0.3$ for FactSelfCheck and $0.75$ for SelfCheckGPT).
Subsequently, we evaluated the factuality of the corrected responses using the LLM-as-judge approach Zheng et al. (2023) using Prompt 10. For each corrected sentence, we provided the external biography from Wikipedia. We instructed LLM-as-judge to return ’yes’ if the source supported the sentence, ’no’ if it was not, or ’refused’ if the LLM declined to correct the sentence (e.g., due to insufficient knowledge). We then categorized the responses into three labels: ’factual’, ’non-factual’, ’refused’.
As mentioned in Section 4.2, we utilized GPT-4o for this experiment instead of Llama-3.1-70B-Instruct (used in detection). This choice was motivated by the challenging nature of the correction task – the model needs to correct hallucinations using only its internal knowledge, without access to external references. While the model knows hallucinated parts, it must rely on its knowledge to determine the correct information.
While the described correction method is not our main contribution, we used it to study the potential benefits of fact-level detection. Although the correction method employed here may not be the most sophisticated, the key takeaway is the observed difference in performance.
5 Results
This section presents the results of the experiments described in the previous section. The results of the fact-level evaluation are given in Appendix B.
5.1 Evaluation of Sentence-Level and Passage-Level Hallucination Detection
Method Agg. AUC-PR
Sentence-level methods
SCGPT (Prompt) - 93.60
SCGPT (NLI) - 92.50
SCGPT (Unigram-max) - 85.63
SCGPT (QA) - 84.26
SCGPT (BERTScore) - 81.96
RandomSentence - 72.96
Fact-level methods (ours)
FSC-Text max 92.45
FSC-KG (LLM-based) max 91.82
FSC-Text mean 91.01
FSC-KG (LLM-based) mean 90.24
FSC-KG (Freq.-based) max 88.48
FSC-KG (Freq.-based) mean 88.25
RandomFact mean 74.22
Table 1: Results on the sentence-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores.
Method Agg. Pear. Spear.
Sentence-level methods
SCGPT (Prompt) - 83.12 78.65
SCGPT (NLI) - 74.14 73.78
SCGPT (Unigram-max) - 64.71 64.91
SCGPT (QA) - 61.07 59.29
SCGPT (BERTScore) - 58.18 55.90
Fact-level methods (ours)
FSC-Text max 79.69 76.68
FSC-KG (LLM-based) max 77.30 75.37
FSC-Text mean 77.03 75.89
FSC-KG (LLM-based) mean 73.92 72.28
FSC-KG (Freq.-based) mean 71.50 70.45
FSC-KG (Freq.-based) max 64.05 66.09
Table 2: Results on the passage-level hallucination detection task. Comparison of sentence-level and fact-level methods based on Pearson and Spearman correlation scores. Results are sorted by the Pearson correlation score. SCGPT stands for SelfCheckGPT, FSC represents FactSelfCheck, and Agg denotes the aggregation method used for calculating sentence-level scores. As mentioned in Section 4, passage-level scores were calculated using the mean of sentence-level scores.
Tables 1 and 2 present a comparative analysis of our method against SelfCheckGPT (SCGPT). In the sentence-level evaluation, FactSelfCheck-Text utilizing $max$ as an aggregation function achieves an AUC-PR score of $92.45$. It demonstrates that our approach is comparable in performance to the leading SCGPT variants – Prompt ($93.60$) and NLI ($92.50$). Notably, while our method operates at a more granular level, it maintains competitive performance with a marginal decrease of $1.2\%$ compared to the best SCGPT. Figure 3 illustrates the precision-recall trade-off across various detection thresholds.
In the passage-level evaluation, our method achieves Pearson and Spearman correlation coefficients of $79.69$ and $76.68$, respectively. Although these results are lower than SCGPT (Prompt) ($83.12$ and $78.65$), they demonstrate the effective scalability of the fact-level detection to the sentence-level and passage-level detections.
Among the FactSelfCheck (FSC) variants, FSC-Text, which performs the direct comparison between facts and samples, consistently outperforms FSC-KG, which relies on knowledge graph comparisons. FSC-Text offers computational advantages by eliminating the knowledge graph extraction step. However, FSC-KG may be more computationally efficient for processing longer samples with a lower fact density because of reduced token usage. Both variants perform better than the Frequency-based FSC, which employs a simple fact occurrence counting mechanism. This performance differential highlights the benefits of incorporating semantic matching and knowledge graph reasoning in hallucination detection. Nevertheless, the frequency-based method’s computational efficiency may make it suitable for applications where its reduced accuracy is acceptable. Regarding aggregation functions, $max$ consistently yields better results than $mean$, as it is more sensitive to the presence of hallucinated facts in the text.
An interesting side observation is that our reproduced SCGPT on Llama-3.1-70B-Instruct marginally surpassed the original implementation using GPT-3.5-turbo ($93.60$ vs $93.42$  [^5] The scores on GPT-3.5-turbo were obtained from the original paper Manakul et al. (2023).). Furthermore, for passage-level detection, Llama achieved a better Pearson correlation coefficient than the GPT-3.5-turbo ($83.12$ vs. $78.32$ 5), while the Spearman correlation coefficient remained at a similar level ($78.65$ vs. $78.30$ 5). This indicates that contemporary open-source models can exceed the performance of previous closed models.
5.2 Effect of Sample Size on Detection Performance
Figure 2: Impact of the number of samples on sentence-level detection task.
Figures 2 and 4 illustrate the relationship between the number of samples and detection performance for both FactSelfCheck variants and SelfCheckGPT (Prompt). The results show that FactSelfCheck exhibits similar behavior to SelfCheckGPT regarding sample requirements. The performance improves dramatically with up to $5$ samples, after which the improvement curve flattens. While additional samples continue to yield benefits, these improvements become incremental, with modest gains observed up to $20$ samples.
5.3 Role of Fact-Level Detection in Hallucination Correction
Level Factual $\uparrow$ Non-Factual $\downarrow$ Refused
baseline 0.23 0.74 0.04
sentence 0.25 (+8%) 0.71 (-4%) 0.05 (+30%)
fact 0.31 (+35%) 0.64 (-12%) 0.05 (+30%)
Table 3: Effectiveness of hallucination correction by providing detected hallucinations at sentence-level, fact-level, and a baseline (without providing any hallucinations). The table presents the proportions of factual, non-factual sentences, and refused corrections. Percentages in parentheses indicate the relative change compared to the baseline. Arrows $\uparrow$ and $\downarrow$ denote whether a higher or lower value is better.
Table 3 presents the results of our hallucination correction experiment. The fact-level approach shows substantial improvements over the baseline and sentence-level methods. We observed a $35\%$ increase in factual content and $12\%$ reduction in non-factual content compared to the baseline. In contrast, the sentence-level detection achieves only modest improvements of $8\%$ and $4\%$, respectively, indicating that pointing out hallucinations at the fact level enables more effective corrections and underscoring the importance of our study and contributions.
The overall rate of refusals remains low, increasing only marginally from $0.04$ in the baseline to $0.05$ with fact-level and sentence-level detection. We hypothesize that the model becomes more cautious with provided information about hallucinations and more likely to know its limitations.
6 Conclusion
Our study introduces FactSelfCheck, a novel approach to hallucination detection that operates at the fact level rather than the sentence level or passage level. The evaluation results demonstrate that this granular approach achieves competitive performance with methods like SelfCheckGPT, with the AUC-PR score of $92.45$ on sentence-level detection compared to SelfCheckGPT (Prompt)’s $93.60$, while offering more detailed insights into the nature of hallucinations.
Our fact-level approach improves hallucination correction. Providing the correction system with a list of incorrect facts rather than incorrect sentences allows for more effective corrections. Compared to baseline, we achieved $35\%$ increase in factual content and $12\%$ reduction in non-factual content. In contrast, the sentence-level approach achieved a $8\%$ increase and $4\%$ reduction, respectively. It underscores the importance of fine-grained hallucination detection in enhancing the reliability and trustworthiness of systems that use LLMs, particularly in applications requiring high factual accuracy.
Limitations
Our study faces two primary limitations. First, we are constrained by the availability of suitable datasets. Currently, the WikiBio GPT-3 Hallucination Dataset, to the best of our knowledge, is the only dataset designed for evaluating sampling-based methods, and it only provides sentence-level and passage-level annotations. It forced us to evaluate our method through aggregation rather than directly assessing our fact-level detection capabilities. Second, while the granular approach of FactSelfCheck justifies its increased complexity, the multiple LLM-based steps make it more computationally intensive compared to more straightforward methods like SelfCheckGPT.
Several promising directions could address these limitations. An important step would be creating new datasets with fact-level hallucination annotations, enabling direct evaluation of our method’s core capabilities. Additionally, we see significant potential for improving computational efficiency. Our current prompt engineering was largely empirical and not optimized for token usage. Future work could focus on reducing prompt lengths and merging steps, such as simultaneously assessing support for multiple facts or merging the KG extraction steps. Finally, incorporating external knowledge graphs could enhance fact verification while potentially reducing our reliance on stochastic sampling.
Ethical Considerations
Like all machine learning methods, FactSelfCheck can produce false positives and false negatives. Therefore, it should not completely replace human verification of factual correctness in LLM responses. Additionally, since the fact-level annotations (Appendix B) are generated by an LLM, there remains a risk of fact mislabeling.
Acknowledgments
This work was funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, http://ominoproject.eu/). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. It was also co-financed with funds from the Polish Ministry of Education and Science under the programme entitled International Co-Financed Projects, grant no. 573977. This work was co-funded by the National Science Centre, Poland under CHIST-ERA Open & Re-usable Research Data & Software (grant number 2022/04/Y/ST6/00183).
Appendix A Additional Figures for Results
Figure 3 depicts the precision-recall curve for sentence-level detection (see Section 5.1). Figure 4 shows the impact of the number of samples on the passage-level detection task (see Section 5.2).
Figure 3: Precision-recall curve for the sentence-level detection task.
Figure 4: Impact of the number of samples on passage-level detection task.
Appendix B Fact-Level Detection Evaluation
While the evaluation in Section 5.1 focused on the sentence-level and passage-level detection due to dataset constraints, we present a complementary evaluation at the fact level here. Since this evaluation is not based on human annotations, we have placed it in the appendix.
Evaluation Data
To enable fact-level evaluation, we annotated facts from the extracted response knowledge graph ($KG_{p}$) using the LLM-as-judge approach Zheng et al. (2023). This approach annotated whether each fact is supported by the external Wikipedia biography (Prompt 11). As a result, we obtained $5488$ binary annotated facts.
Experimental Setup
We evaluated all variants of FactSelfCheck using the fact-level score $\mathcal{H}_{\text{fact}}$. For comparison with SelfCheckGPT, which provides only sentence-level granularity, we derived fact-level scores by averaging the sentence-level scores across all sentences containing each fact.
Results
Table 4 presents the comparative results. FactSelfCheck-Text demonstrates superior performance with an AUC-PR score of $93.41$, followed by FactSelfCheck-KG (LLM-based) at $92.25$. The frequency-based method achieves $87.99$, while SelfCheckGPT (Prompt) scores $86.18$. These results demonstrate the effectiveness of our method in detecting hallucinations at the fact level. Furthermore, the lower performance of averaging the sentence-level scores highlights the importance of designing fact-level methods and validating our approach.
Method AUC-PR
FactSelfCheck-Text 93.41
FactSelfCheck-KG (LLM-based) 92.25
FactSelfCheck-KG (Freq.-based) 87.99
SelfCheckGPT (Prompt) 86.18
RandomFact 65.79
Table 4: Results on the fact-level hallucination detection task. Comparison of sentence-level and fact-level methods based on AUC-PR scores.
Appendix C Passage-level Score: Size of Knowledge Graphs Extracted from Samples
As a side study, we examined the effectiveness of scoring hallucinations at the passage level based on the size of knowledge graphs extracted from samples.
Score Definition
We propose using the average size of restricted knowledge graphs extracted from samples as a passage-level score. Restriction involves keeping only facts where all entities and relations are present in $\mathcal{E}_{S}$ or $\mathcal{R}_{S}$, respectively. This results in $KG_{s}^{\prime}$ defined as:
[…]
$h\in\mathcal{E}_{S}\land r\in\mathcal{R}_{S}\land t\in\mathcal{E}_{S}\}$
The passage-level score is then calculated as:
$\mathcal{H}_{\text{passage}}(p)=\frac{1}{|S|}\sum_{s\in S}-|KG_{s}^{\prime}|$ (14)
The rationale behind this score is that the size of $KG_{s}^{\prime}$ reflects how well the samples align with the original response $p$. A smaller $KG_{s}^{\prime}$ indicates that the sample contains facts than those in the original response, suggesting potential hallucinations. In distinction from the previously defined scores, this metric operates outside the standard $[0,1]$ range.
Results
The method achieved Pearson and Spearman correlation scores of $72.35$ and $71.18$, respectively. When compared to the main results in Table 2, this approach shows lower correlation scores than LLM-based methods but outperforms the frequency-based method. It is important to note that while it is the passage-level score, frequency-based methods operate at the fact level. Nevertheless, this approach offers the most computationally efficient way to calculate passage-level scores among tested methods utilizing knowledge graphs, as it does not require any LLM calls during factual consistency assessment.
Appendix D Case Study of FactSelfCheck vs SelfCheckGPT
Table 5 and Figure 5 present a comparative case study of predictions made by FactSelfCheck and SelfCheckGPT. The table contains an external Wikipedia biography, sentences from the response, facts extracted from the response, and the predictions of both methods. This comparison demonstrates that fact-level detection provides more detailed information about the factuality of the response. We observe that LLM did not hallucinate all facts, as some are consistent with the external Wikipedia biography. However, when using sentence-level detection, we cannot distinguish between correct and hallucinated facts – all sentences are predicted as hallucinated.
Appendix E SelfCheckGPT with Enhanced Prompt
To ensure a fair comparison between FactSelfCheck and SelfCheckGPT, we conducted an additional experiment using an enhanced prompt for SelfCheckGPT. The original SelfCheckGPT prompt is relatively simple, while our FactSelfCheck prompts are more elaborate, directly allowing reasoning and inference of new facts, and providing examples. These characteristics could potentially increase the performance of methods. We designed an alternative prompt for SelfCheckGPT, presented in Prompt 12, that incorporates these features, making it similar to our FactSelfCheck prompts.
For sentence-level detection, the enhanced prompt for SelfCheckGPT achieved an AUC-PR score of $93.38$, slightly lower than the original prompt’s $93.60$. This minimal difference indicates that SelfCheckGPT’s performance is not significantly affected by prompt design in our experimental setting. In fact, the enhanced prompt even lowered the performance rather than increasing it, despite its more sophisticated design. For passage-level detection, we observed similar results, with the enhanced prompt achieving Pearson and Spearman correlation scores of $82.36$ and $76.71$, respectively, compared to the original prompt’s $83.12$ and $78.65$.
Due to these findings, we chose to use the original prompt for SelfCheckGPT in our main experiments. These consistent results confirm that our comparison between methods is fair, as the enhanced prompt did not improve the performance of SelfCheckGPT.
Appendix F FactSelfCheck Prompts
In this section, we present the prompts used in FactSelfCheck. These include prompts for the entity extraction (Prompt 1), the relation extraction (Prompt 2), and the knowledge graph extraction from a sentence (Prompt 3) or a sample (Prompt 4). Additionally, we provide prompts for assessing factual consistency in FactSelfCheck-KG (Prompt 5) and FactSelfCheck-Text (Prompt 6).
Appendix G Correction Prompts
The following prompts are instructions for correcting hallucinated responses generated by the LLM. The prompts are for the baseline (Prompt 7), sentence-level (Prompt 8), and fact-level (Prompt 9) correction. The prompts take several inputs: {generated_sentences} containing the response as a numbered list of sentences, {incorrect_sentences} listing the hallucinated sentences with their corresponding numbers from the generated text, and {incorrect_facts} providing a formatted list of hallucinated facts organized by sentence. The {format} field specifies the output format as an empty numbered list with a length equal to the number of sentences in the {generated_sentences} field.
External Wikipedia Bio
Kenan Hasagić (born 1 February 1980) is a Bosnian football goalkeeper who plays for Balıkesirspor. His football career began in his hometown with FK Rudar. At the age of 16, he made his debut in a first division match. He was the most promising goalkeeper in Bosnia and Herzegovina; he played for youth selections and was later transferred to Austrian side Vorwärts Steyr. After that, he was a member of Altay SK in Turkey but didn’t see much first team football. He went back to Bosnia and played for Bosna Visoko. In 2003, he signed a contract with FK Željezničar. Here he found good form and even became first choice goalkeeper for the Bosnian national team. In the 2004–05 season, he moved to Turkey once again where he signed for Turkish Süper Lig side Gaziantepspor. He made his debut for the national team on 12 February 2003 in a game between Wales and Bosnia and Herzegovina which ended in a 2–2 draw.
Sentence 1 (SelfCheckGPT: 1.0)
Kenan Hasagić (born 28 April 1988) is a Bosnian professional footballer who plays as a midfielder for Bosnian Premier League club Željezničar.
Fact FactSelfCheck
(’Kenan Hasagić’, ’DATE OF BIRTH’, ’April 28, 1988’) 1.00
(’Kenan Hasagić’, ’NATIONALITY’, ’Bosnian’) 0.00
(’Kenan Hasagić’, ’OCCUPATION’, ’professional footballer’) 0.20
(’Kenan Hasagić’, ’POSITION PLAYED’, ’midfielder’) 0.55
(’Kenan Hasagić’, ’LEAGUE’, ’Bosnian Premier League’) 0.45
(’Kenan Hasagić’, ’CURRENT CLUB’, ’Željezničar’) 0.85
Sentence 2 (SelfCheckGPT: 1.0)
Hasagić started his career at his hometown club Željezničar, where he made his professional debut in 2006.
Fact FactSelfCheck
(’Kenan Hasagić’, ’CURRENT CLUB’, ’Željezničar’) 0.85
(’Kenan Hasagić’, ’PROFESSIONAL DEBUT’, ’2006’) 0.90
(’Kenan Hasagić’, ’CURRENT CLUB’, ’Željezničar’) 0.85
Sentence 3 (SelfCheckGPT: 1.0)
He has since gone on to make over 200 appearances for the club, winning the Bosnian Premier League title in 2008 and the Bosnian Cup in 2009.
Fact FactSelfCheck
(’Kenan Hasagić’, ’TEAM APPEARANCES’, ’Željezničar’) 0.60
(’Željezničar’, ’CHAMPIONSHIP WON’, ’Bosnian Premier League title’) 0.90
(’Željezničar’, ’YEAR OF CHAMPIONSHIP’, ’2008’) 1.00
(’Željezničar’, ’CHAMPIONSHIP WON’, ’Bosnian Cup’) 0.95
(’Željezničar’, ’YEAR OF CHAMPIONSHIP’, ’2009’) 1.00
Sentence 4 (SelfCheckGPT: 0.9)
He has also represented Bosnia and Herzegovina at international level, making his debut in 2011.
Fact FactSelfCheck
(’Kenan Hasagić’, ’COUNTRY REPRESENTED’, ’Bosnia and Herzegovina’) 0.05
(’Kenan Hasagić’, ’INTERNATIONAL DEBUT’, ’2011’) 0.85
Table 5: Comparison of fact-level FactSelfCheck with sentence-level SelfCheckGPT. An external Wikipedia biography is provided to analyse the correctness of the methods. The red value indicates hallucinations, and the green value indicates factual correctness. The facts were classified using a threshold of $0.4$ utilizing FactSelfCheck, and the sentences were classified using a threshold of $0.75$ with SelfCheckGPT. These thresholds achieved the highest F1-scores in fact-level (Appendix B) and sentence-level (Section 5.1) evaluation, respectively.
Figure 5: Example of a knowledge graph extracted from a response. Edge width represents the hallucination score for each fact, with red edges indicating hallucinated facts and green edges indicating correct facts. Facts were classified using a threshold of $0.4$, which achieved the highest F1-score in the fact-level evaluation (Appendix B).
================================ System Message ================================
You are a top-tier algorithm designed for extracting entities in structured formats.
These entities will be used to build a knowledge graph.
Your task is to identify the entities requested with the user prompt from a given text.
You must generate the output in a JSON format containing a list with JSON objects.
Each object should have the key: "entities".
The "entities" key must contain a list of strings, where each string is an extracted entity.
Attempt to extract as many entities as you can.
It is very important to cover all the entities in the text!
Entities can be in many types, such as people, organizations, locations, events, positions, dates, roles, titles, objects, and more.
If an entity, such as "John Doe", is mentioned multiple times in the text but is referred to by different names or pronouns (e.g., "Joe", "he"), always use the most complete identifier for that entity.
IMPORTANT NOTES:
- Dont add any explanation and text.
- It is very important to cover all the entities in the text!
- Remember to cover dates.
- Remember to cover events eg. publication of book, employment, graduation, marriage.
- All numbers, dates has to be in string format. eg. "2023", "1980s".
================================ Human Message =================================
Below are a number of examples of text and their extracted entities.
[{’text’: "Donald John Trump (born June 14, 1946) is an American politician, media personality, and businessman who served as the 45th president of the United States from 2017 to 2021. Having won the 2024 presidential election as the nominee of the Republican Party, he is the president-elect and will be inaugurated as the 47th president on January 20, 2025. Trump graduated with a bachelors degree in economics from the University of Pennsylvania in 1968.", entities’: [’Donald John Trump’, American politician’, media personality’, businessman’, ’45th president’, United States’, ’2017’, ’2021’, ’2024 presidential election’, Republican Party’, president-elect’, ’47th president’, January 20, 2025’, University of Pennsylvania’, graduation’, bachelor’, degree’, economics’, ’1968’]}]
For the following text, extract entities as in the provided example.
{format_instructions}
Text: {input}
Prompt 1: Instructions for the LLM to extract entities $\mathcal{E}_{p}$ from the response ({input}). The {format_instructions} field contains formatting instructions for JSON output, constructed by the LangChain library based on the expected output schema.
================================ System Message ================================
You are a top-tier algorithm designed for extracting relationship types in structured formats.
These relationship types will be used to build a knowledge graph in the future.
Your task is to identify the relationship types requested with the user prompt from a given text and list of entities.
You must generate the output in a JSON format containing a list with JSON objects.
Each object should have the key: "relationship_types".
The "relationship_types" key must contain a list of strings, where each string is an extracted relationship type.
Attempt to extract as many relationship types as you can.
It is very important to cover all the relationship types in the text!
Ensure the relationships accurately describe the interaction or connection between entities.
Relationship types should be simple and concise, but also specific and informative, to cover all the relationships between entities in the text.
If it is possible, use the WikiData relationship types. If not, use the most appropriate relationship type.
IMPORTANT NOTES:
- Dont add any explanation and text.
- It is very important to cover all the relationship types in the text!
- Omit long and complex relationship types.
- Use double quotes, not single quotes for json format.
================================ Human Message =================================
Below are a number of examples of text and their extracted relationship types.
[{’text’: "Donald John Trump (born June 14, 1946) is an American politician, media personality, and businessman who served as the 45th president of the United States from 2017 to 2021. Having won the 2024 presidential election as the nominee of the Republican Party, he is the president-elect and will be inaugurated as the 47th president on January 20, 2025. Trump graduated with a bachelors degree in economics from the University of Pennsylvania in 1968.", entities’: [’Donald John Trump’, American politician’, media personality’, businessman’, ’45th president’, United States’, ’2017’, ’2021’, ’2024 presidential election’, Republican Party’, president-elect’, ’47th president’, January 20, 2025’, University of Pennsylvania’, graduation’, bachelor’, degree’, economics’, ’1968’], relationship_types’: [’DATE OF BIRTH’, NATIONALITY’, OCCUPATION’, POSITION HELD’, TERM OF OFFICE’, MEMBER OF’, ELECTED AS’, NOMINEE OF’, FUTURE POSITION’, DATE OF START’, DATE OF END’, EDUCATED AT’, DEGREE IN’, DEGREE FROM’, DEGREE TYPE’, DATE OF GRADUATION’]}]
For the following text and entities, extract relationship types as in the provided example.
{format_instructions}
Text: {input}
Entities: {entities}
Prompt 2: Instructions for the LLM to extract relations $\mathcal{R}_{p}$ from the response ({input}). The {entities} field contains entities extracted using Prompt 1. The {format_instructions} field specifies JSON output formatting requirements constructed by the LangChain library based on the expected schema.
================================ System Message ================================
You are a top-tier algorithm designed for extracting information in CSV format to build a knowledge graph.
In input, you will receive: sentence, full text, list of entities, and list of relationships.
Your task is to identify facts (triples) consisting of a subject, relation, and object from a given sentence.
Do not generate relationships that are defined in the full text but not in the sentence.
[…]
The first line should be the header: subject;relation;object
Each subsequent line should contain one triple with the following columns:
- subject: the text of the extracted entity which is the subject of the relation
- relation: the type of relation between the subject and object
- object: the text of the extracted entity which is the object of the relation
The extracted entities must be from the list provided by the user.
The relations must be from the list of allowed relation types provided by the user.
Attempt to extract as many facts / triples as you can.
Entities can be in many types, such as people, organizations, locations, dates, roles, titles, objects, and more.
Maintain Entity Consistency: When extracting entities, its vital to ensure consistency.
If an entity, such as "John Doe", is mentioned multiple times in the text but is referred to by different names or pronouns (e.g., "Joe", "he"), always use the most complete identifier for that entity.
The knowledge graph should be coherent and easily understandable, so maintaining consistency in entity references is crucial.
The knowledge graph should be complete, so it should cover all the facts in the sentence.
IMPORTANT NOTES:
- Dont add any explanation and text.
- It is very important to cover all triples (facts) in the sentence!
- Full text is only for understanding the context. Do not add any triples that are not in the sentence.
[…]
- Do not include quotes around the values unless they contain semicolons
================================ Human Message =================================
For the following sentence, full text, entities, and relation types, extract facts / triples as in the provided sentence.
Your response should be a list of semicolon separated values with header, eg:
subject;relation;object
Donald John Trump;DEGREE FROM;University of Pennsylvania
Donald John Trump;DEGREE IN;economics
Below are a number of examples of text and their extracted facts.
[…]
Use the following entities, dont use other entity that is not defined below:
# ALLOWED ENTITIES:
{allowed_nodes}
Use the following relation types, dont use other relation that is not defined below:
# ALLOWED RELATION TYPES:
{allowed_relationships}
Full text: {input_text}
Sentence: {input_sentence}
Prompt 3: Instructions for the LLM to extract knowledge graph $KG_{u}$ from a sentence ({input_sentence}). We also provide a full response ({input_text}) to give LLM a context. The {allowed_nodes} field contains the list of allowed entities. The {allowed_relationships} field contains the list of allowed relations. The LLM is expected to return a list of triples in the CSV format.
================================ System Message ================================
You are a top-tier algorithm designed for extracting information in CSV format to build a knowledge graph.
Your task is to identify facts (triples) consisting of a subject, relation, and object from a given text.
[…]
The first line should be the header: subject;relation;object
Each subsequent line should contain one triple with the following columns:
- subject: the text of the extracted entity which is the subject of the relation
- relation: the type of relation between the subject and object
- object: the text of the extracted entity which is the object of the relation
The extracted entities must be from the list provided by the user.
The relations must be from the list of allowed relation types provided by the user.
Attempt to extract as many facts / triples as you can.
Entities can be in many types, such as people, organizations, locations, dates, roles, titles, objects, and more.
Maintain Entity Consistency: When extracting entities, its vital to ensure consistency.
If an entity, such as "John Doe", is mentioned multiple times in the text but is referred to by different names or pronouns (e.g., "Joe", "he"), always use the most complete identifier for that entity.
The knowledge graph should be coherent and easily understandable, so maintaining consistency in entity references is crucial.
The knowledge graph should be complete, so it should cover all the facts in the text.
IMPORTANT NOTES:
- Dont add any explanation and text.
- It is very important to cover all triples (facts) in the text!
[…]
- Do not include quotes around the values unless they contain semicolons
================================ Human Message =================================
For the following text, allowed entities and allowed relation types, extract facts / triples as in the provided example.
Your response should be a list of semicolon separated values with header, eg:
subject;relation;object
Donald John Trump;DATE OF BIRTH;June 14, 1946
Donald John Trump;NATIONALITY;American politician
Below are a number of examples of text and their extracted facts.
[…]
Use the following entities, dont use other entity that is not defined below:
# ALLOWED ENTITIES:
{allowed_nodes}
Use the following relation types, dont use other relation that is not defined below:
# ALLOWED RELATION TYPES:
{allowed_relationships}
Text: {input}
Prompt 4: Instructions for the LLM to extract a knowledge graph $KG_{s}$ from a sample ({input}). The {allowed_nodes} field contains the list of allowed entities. The {allowed_relationships} field contains the list of allowed relations. The LLM is expected to return a list of triples in the CSV format.
================================ System Message ================================
You are a top-tier algorithm designed for verification whether fact is supported by the provided knowledge graph.
Your task is to answer whether the given fact is supported by the knowledge graph.
Return yes if the fact is supported, no otherwise.
You can reason over the knowledge graph and the fact to answer.
You can infer the answer from the knowledge graph, e.g. by inferring new facts from the knowledge graph.
Do not add new facts other than by inference from the knowledge graph.
Fact is a triple of the form ‘(subject, predicate, object)‘. Knowledge graph is a list of facts.
You must return a single word, either yes or no‘.
You must not return any other text or explanation.
================================ Human Message =================================
Below are a number of examples of facts, knowledge graph and their verification results.
- Fact: (Donald Trump, president, United States)
Knowledge graph: [(Donald Trump, president, United States), (Donald Trump, born, 1946), (Donald Trump, graduated, University of Pennsylvania)]
Is the fact supported by the knowledge graph?: yes
- Fact: (Joe Biden, born, 1942)
Knowledge graph: [(Joe Biden, proffesion, politician), (Joe Biden, president, United States), (Joe Biden, graduated, University of Delaware)]
Is the fact supported by the knowledge graph?: no
- Fact: (Michael Jordan, played, basketball)
Knowledge graph: [(Michael Jordan, CEO, PepsiCo), (Michael Jordan, born, 1936), (Michael Jordan, graduated, Yale University)]
Is the fact supported by the knowledge graph?: no
- Fact: (Kobe Bryant, played in, NBA)
Knowledge graph: [(Kobe Bryant, born, 1978), (Kobe Bryant, played in, Los Angeles Lakers)]
note: Los Angeles Lakers is a part of NBA
Is the fact supported by the knowledge graph?: yes
- Fact: (Bryant, profession, basketball player)
Knowledge graph: [(Kobe Bryant, born, 1978), (Kobe Bryant, played in, Los Angeles Lakers)]
note: In this context, Bryant is Kobe Bryant, and Los Angeles Lakers is a basketball team, so Kobe Bryant is a basketball player
Is the fact supported by the knowledge graph?: yes
For the following fact and knowledge graph, verify if the fact is supported by the knowledge graph.
Fact: {input}
Knowledge graph: {knowledge_graph}
Is the fact supported by the knowledge graph?
Prompt 5: FactSelfCheck-KG: Instructions for the LLM to assess support of fact ({input}) by a knowledge graph ({knowledge_graph}). The LLM is expected to return ’yes’ if the fact is supported by the knowledge graph, ’no’ otherwise.
================================ System Message ================================
You are a top-tier algorithm designed for verification whether fact is supported by the provided context.
Your task is to answer whether the given fact is supported by the context.
Return yes if the fact is supported, no otherwise.
You can reason over the context and the fact to answer.
You can infer the answer from the context, e.g. by inferring new facts from the context.
Do not add new facts other than by inference from the context.
Fact is a triple of the form ‘(subject, predicate, object)‘. Context is a text.
You must return a single word, either yes or no‘.
You must not return any other text or explanation.
================================ Human Message =================================
Below are a number of examples of facts, context and their verification results.
- Fact: (Donald Trump, president, United States)
Context: Donald Trump was president of the United States. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
Is the fact supported by the context?: yes
- Fact: (Joe Biden, born, 1942)
Context: Joe Biden is a politician and president of the United States. He graduated from the University of Delaware.
Is the fact supported by the context?: no
- Fact: (Michael Jordan, played, basketball)
Context: Michael Jordan is the CEO of PepsiCo. He was born in 1936 and graduated from Yale University.
Is the fact supported by the context?: no
- Fact: (Kobe Bryant, played in, NBA)
Context: Kobe Bryant was born in 1978 and played for the Los Angeles Lakers.
note: Los Angeles Lakers is a part of NBA
Is the fact supported by the context?: yes
- Fact: (Bryant, profession, basketball player)
Context: Kobe Bryant was born in 1978 and played for the Los Angeles Lakers.
note: In this context, Bryant is Kobe Bryant, and Los Angeles Lakers is a basketball team, so Kobe Bryant is a basketball player
Is the fact supported by the context?: yes
For the following fact and context, verify if the fact is supported by the context.
Fact: {input}
Context: {context}
Is the fact supported by the context?
Prompt 6: FactSelfCheck-Text: Instructions for the LLM to assess support of fact ({input}) by a sample ({context}). The LLM is expected to return ’yes’ if the fact is supported by the sample, ’no’ otherwise.
================================ System Message ================================
You are a top-tier algorithm designed for correcting sentences from the generated text generated by the LLM.
You will be given a original prompt, a generated text in form of a list of sentences.
We dont know whether the given sentences are correct or incorrect.
You need to correct the incorrect sentences to make it more factual and coherent.
If the sentence is correct, you should return the original sentence.
If you dont know how to correct the sentence, you should return the original sentence.
Put the corrected sentences in the following format, each sentence separated by a new line:
1. Corrected sentence 1.
2. Corrected sentence 2.
You must return only the corrected or original sentences, not return any other text or explanation.
Important: The number of corrected sentences should be the same as the number of sentences in the generated text.
================================ Human Message =================================
Below are a number of examples of original prompt, list of sentences from the generated text.
- Original prompt: ‘‘‘Donald Trump is‘‘‘
Generated sentences:
1. Donald Trump is the president of Canada.
2. He was born in 1946 and graduated from the University of Pennsylvania.
Corrected sentences:
1. Donald Trump is the president of the United States.
2. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
- Original prompt: ‘‘‘Kobe Bryant is‘‘‘
Generated sentences:
1. Kobe Bryant (February 17, 1963 - January 26, 2020) was an American professional basketball player.
2. Bryant married Vanessa Laine in 2001.
Corrected sentences:
1. Kobe Bryant (August 23, 1978 - January 26, 2020) was an American professional basketball player.
2. Kobe Bryant married Vanessa Laine in 2001.
For the following original prompt, generated text, and incorrect sentences, correct the generated text.
Original prompt: ‘‘‘{original_prompt}‘‘‘
Generated sentences:
{generated_sentences}
Answer in format:
{format}
Corrected sentences:
Prompt 7: Instructions for the LLM to correct a hallucinated response without pointing out hallucinations. For a description of input fields { }, see Appendix G.
================================ System Message ================================
You are a top-tier algorithm designed for correcting sentences from the generated text generated by the LLM.
You will be given a original prompt, sentences from the generated text, and a list of incorrect sentences from the generated text.
Empty list means that no incorrect sentences were found.
You need to correct the sentences to make it more factual and coherent.
If the sentence is correct, you should return the original sentence.
If you dont know how to correct the sentence, you should return the original sentence.
Put the corrected sentences in the following format, each sentence separated by a new line:
1. Corrected sentence 1.
2. Corrected sentence 2.
3. Corrected sentence 3.
You must return only the corrected or original sentences, not return any other text or explanation.
Important: The number of corrected sentences should be the same as the number of sentences in the generated text.
================================ Human Message =================================
Below are a number of examples of original prompt, sentences from the generated text, and incorrect sentences.
- Original prompt: ‘‘‘Donald Trump is‘‘‘
Generated sentences:
1. Donald Trump is the president of Canada.
2. He was born in 1946 and graduated from the University of Pennsylvania.
Incorrect sentences:
1. Donald Trump is the president of Canada.
Corrected sentences:
1. Donald Trump is the president of the United States.
2. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
- Original prompt: ‘‘‘Kobe Bryant is‘‘‘
Generated sentences:
1. Kobe Bryant (February 17, 1963 - January 26, 2020) was an American professional basketball player.
2. Bryant married Vanessa Laine in 2001.
Incorrect sentences:
1. Kobe Bryant (February 17, 1963 - January 26, 2020) was an American professional basketball player.
Corrected sentences:
1. Kobe Bryant (August 23, 1978 - January 26, 2020) was an American professional basketball player.
2. Kobe Bryant married Vanessa Laine in 2001.
For the following original prompt, sentences from the generated text, and incorrect sentences, correct the generated text.
Original prompt: ‘‘‘{original_prompt}‘‘‘
Generated sentences:
{generated_sentences}
Incorrect sentences:
‘‘‘{incorrect_sentences}‘‘‘
Answer in format:
{format}
Corrected sentences:
Prompt 8: Instructions for the LLM to correct a hallucinated response by pointing out hallucinated sentences. For a description of input fields { }, see Appendix G.
================================ System Message ================================
You are a top-tier algorithm designed for correcting sentences from the generated text generated by the LLM.
You will be given a original prompt, sentences from the generated text, and list of incorrect facts for each sentence.
Empty list means that no incorrect facts were found.
Fact is a triple of the form ‘(subject, predicate, object)‘, just like in the knowledge graph.
You need to correct the sentences to make it more factual and coherent.
Use the provided incorrect facts to correct the sentences, by locating the incorrect facts in the sentences and replacing them or removing them.
Put the corrected sentences in the following format, each sentence separated by a new line:
1. Corrected sentence 1.
2. Corrected sentence 2.
3. Corrected sentence 3.
You must return only the corrected or original sentences, not return any other text or explanation.
Important: The number of corrected sentences should be the same as the number of sentences in the generated text.
================================ Human Message =================================
Below are a number of examples of original prompt, sentences from the generated text, and incorrect sentences from the generated text.
- Original prompt: ‘‘‘Donald Trump is‘‘‘
Generated sentences:
1. Donald Trump is the president of Canada.
2. He was born in 1946 and graduated from the University of Pennsylvania.
Incorrect facts:
1. [(Donald Trump, president, Canada)]
2. []
Corrected sentences:
1. Donald Trump is the president of the United States.
2. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
- Original prompt: ‘‘‘Kobe Bryant is‘‘‘
Generated sentences:
1. Kobe Bryant (February 17, 1963 - January 26, 2020) was an American professional basketball player.
2. Bryant married Vanessa Laine in 2001.
Incorrect facts:
1. [(Kobe Bryant, born, February 17 1963)]
2. []
Corrected sentences:
1. Kobe Bryant (August 23, 1978 - January 26, 2020) was an American professional basketball player.
2. Kobe Bryant married Vanessa Laine in 2001.
For the following original prompt, sentences from the generated text, and incorrect sentences, correct the generated sentences.
Original prompt: ‘‘‘{original_prompt}‘‘‘
Generated sentences:
{generated_sentences}
Incorrect facts:
{incorrect_facts}
Answer in format:
{format}
Corrected sentences:
Prompt 9: Instructions for the LLM to correct a hallucinated response by pointing out hallucinated facts. For a description of input fields { }, see Appendix G.
================================ System Message ================================
You are a top-tier algorithm designed for verification whether sentence are supported by the source.
You will be given a sentence, full text, and source.
The full text, as well as the sentence, are generated by the LLM.
Evaluate correctness of the sentence, not the full text.
You have to access the factual consistency of the sentence with the source.
If the sentence is fully supported by the source, you should return yes‘.
If some parts of the sentence are not supported by the source, you should return no‘.
If LLM refused to generate the sentence, e.g. LLM didnt have enough information to generate the sentence, you should return refused‘.
You must return a single word, either yes or no or refused‘.
You must not return any other text or explanation.
================================ Human Message =================================
Below are a number of examples of sentence, full text, and source and their verification results.
- Source: Donald Trump was president of the United States. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
Full text: Donald Trump was president of the United States. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
Sentence: Donald Trump was president of the United States.
Is the sentence supported by the source?: yes
- Source: Michael Jordan is the CEO of PepsiCo. He was born in 1936 and graduated from Yale University.
Full text: Michael Jordan was born in 1963. He was a basketball player.
Sentence: He was a basketball player.
Is the sentence supported by the source?: no
- Source: Kobe Bryant was born in 1978 and played for the Los Angeles Lakers.
Full text: Kobe Bryant was born in 1978. He played in the NBA.
Sentence: He played in the NBA.
note: Los Angeles Lakers is a part of NBA
Is the sentence supported by the source?: yes
For the following sentence, full text, and source, verify if the sentence is supported by the source.
Source: {source}
Full text: {full_text}
Sentence: {input}
Is the sentence supported by the source?
Prompt 10: Instructions for the LLM to evaluate factuality of corrected sentences. The prompt takes three inputs: an external Wikipedia biography ({source}), a complete corrected response ({full_text}), and a specific sentence to evaluate ({input}).
================================ System Message ================================
You are a top-tier algorithm designed for annotate whether fact is supported by the source.
You will be given a fact and source.
Fact is a triple of the form ‘(subject, predicate, object)‘, just like in the knowledge graph.
You have to access the factual consistency of the fact with the source.
You must return a single word, either yes or no‘.
You must not return any other text or explanation.
================================ Human Message =================================
Below are a number of examples of fact and source and their verification results.
- Source: Donald Trump was president of the United States. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
Fact: (’Donald Trump’, PRESIDENT’, Canada’)
Is the fact supported by the source?: no
- Source: Michael Jordan is the CEO of PepsiCo. He was born in 1936 and graduated from Yale University.
Fact: (’Michael Jordan’, CEO’, PepsiCo’)
Is the fact supported by the source?: yes
For the following fact and source, verify if the fact is supported by the source.
Source: {source}
Fact: {fact}
Is the fact supported by the source?
Prompt 11: Instructions for the LLM to annotate hallucinations at the fact level. The LLM is asked to determine whether a fact ({fact}) is supported by an external Wikipedia biography ({source}).
================================ System Message ================================
You are a top-tier algorithm designed for verification whether sentence is supported by the provided context.
Your task is to answer whether the given sentence is supported by the context.
Return yes if the sentence is supported, no otherwise.
You can reason over the context and the sentence to answer.
You can infer the answer from the context, e.g. by inferring new facts from the context.
Do not add new facts other than by inference from the context.
You must return a single word, either yes or no‘.
You must not return any other text or explanation.
================================ Human Message =================================
Below are a number of examples of context and sentences, and their verification results.
- Context: Donald Trump was president of the United States. Donald Trump was born in 1946 and graduated from the University of Pennsylvania.
Sentence: Donald Trump served as president of the United States.
Is the sentence supported by the context above? yes
- Context: Joe Biden is a politician and president of the United States. He graduated from the University of Delaware.
Sentence: Joe Biden was born in 1942.
Is the sentence supported by the context above? no
- Context: Michael Jordan is the CEO of PepsiCo. He was born in 1936 and graduated from Yale University.
Sentence: Michael Jordan played basketball.
Is the sentence supported by the context above? no
- Context: Kobe Bryant was born in 1978 and played for the Los Angeles Lakers.
Sentence: Kobe Bryant played in the NBA.
note: Los Angeles Lakers is a part of NBA
Is the sentence supported by the context above? yes
- Context: Kobe Bryant was born in 1978 and played for the Los Angeles Lakers.
Sentence: Bryant was a basketball player.
note: In this context, Bryant is Kobe Bryant, and Los Angeles Lakers is a basketball team, so Kobe Bryant is a basketball player
Is the sentence supported by the context above? yes
For the following context and sentence, verify if the sentence is supported by the context.
Context: {context}
Sentence: {sentence}
Is the sentence supported by the context above?
Prompt 12: Enhanced prompt for SelfCheckGPT.